OpenStudy (anonymous):

Sally said that when you add or subtract positive or negative decimals, you can change the order of the decimals and still get the same answer. Use these decimals to answer the questions: –6.2, 5.71, –3.96, 2.84 Is Sally’s statement true for addition and subtraction of three positive or negative decimals? Write two pairs of equations (one pair for addition, and one pair for subtraction) that support your answer. help please ill fan and medal

OpenStudy (anonymous):

Nikki..........

OpenStudy (anonymous):

huh...

OpenStudy (anonymous):

dun dun duuuuuuun

OpenStudy (anonymous):

lol can you help i have no idea what im doing

OpenStudy (just_one_last_goodbye):

I think it's always true

OpenStudy (just_one_last_goodbye):

-7+(-3)=-10 -3+(-7)=-10 -7-(-3)=-4 -3-(-7)=4 Its sometimes true

OpenStudy (just_one_last_goodbye):

@bruh_its_haley am I right???

OpenStudy (anonymous):

ok thanks

OpenStudy (just_one_last_goodbye):

No prob ^_^